To receive a A in math, tanya must average at least 90 on five tests. Scores on her first four tests were 90, 91, 81, and 92. What is the lowest score that tanya can geton her last test and still receive an A?
Answer by rfer(16322)

(90+91+81+92+x)/5=90
(354+x)/5=90
354+x=90*5
x=450-354
x=96